# From Compression to Expansion: A Layerwise Analysis of In-Context Learning

### v1

@@ C.1 Definition
where $A_{h,\ell}$ represents the attention map of head $h$ at layer $\ell$, and the loss $\mathcal{L}$ is the cross-entropy calculated between the logits of the last token and the ground-truth label. Thus, a saliency map quantifies the importance of internal activations by integrating both attention strength and its gradient-based influence on the model’s outcome.
C.2 Visualization for all layers
In Figure 4, we analyze three layer groups: shallow (average of layers 1-14), intermediate (average of layers 15-16), and deep (layers 17 and above). Figure 17 serves as a supplementary visualization for Figure 4. We show saliency maps of all layers using three demonstrations. In shallow layers, there is more interaction within demonstrations, indicating that the model extracts task information from each demonstration. In deep layers, there is less interaction within demonstrations and more interaction with the last token, indicating that the model uses the accumulated task information to generate output.
@@ C.2 Visualization for all layers
Appendix D Proof of Theorem 5.1
@@
Appendix E Validation of the I.I.D. Assumption under Repitition
In Theorem 5.1, we assume that each demonstration $\bm{h}_{i},i=1,\ldots,K$ is i.i.d. randomly generated from a distribution $\mathcal{H}$ on $\mathbb{R}^{d}$. To validate the importance of this assumption, we design experiments that increase the ICL length by providing additional demonstrations as input. There are two possible extension methods:
@@
As shown in Figure 19 and Figure 19, in distinct mode, as new demonstrations introduce novel information, the accuracy increases and TDNV decreases. However, in repeat mode, since repeated demonstrations add no new task information, the accuracy and TDNV remain largely unchanged. This proves that when the i.i.d. condition is violated, increasing the number of demonstrations does not lead to better performance and compression.
Appendix F Task vector accuracy & Early-exit accuracy
Task vector accuracy[9] refers to how accurately a LLM can perform a task using only a learned representation of the task(task vector), instead of full in-context demonstrations. As shown in Figure 20, to evaluate the task vector accuracy, we conduct the following steps:
@@ Appendix F Task vector accuracy & Early-exit accuracy
where $[]_{:,K+1}$ means the $(K+1)$-th column, $\ell$ is the layer with most compression.
@@
Early-exit accuracy [10, 48] measures a model’s prediction accuracy when using intermediate-layer representations instead of the final layer for making predictions. This metric helps assess how effectively different layers encode the information needed to complete the task. Let $\bm{h}^{(\ell)}\in\mathbb{R}^{d}$ be the hidden state of the final token at layer $\ell$, and let $\mathcal{C}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{V}$ be the last-layer classifier mapping from hidden dimension $d$ to vocabulary size $V$. Then as shown in Figure 21, the prediction is,

### v2

@@ D.1 Saliency Maps
where $A_{h,\ell}$ represents the attention map of head $h$ at layer $\ell$, and the loss $\mathcal{L}$ is the cross-entropy calculated between the logits of the last token and the ground-truth label. Thus, a saliency map quantifies the importance of internal activations by integrating both attention strength and its gradient-based influence on the model’s outcome. In a nutshell, these maps highlight how token interactions evolve across layers.
We show saliency maps of all layers using three demonstrations in Figure˜19. In shallow layers, there is more interaction within demonstrations, indicating that the model extracts task information from each demonstration. In deep layers, there is less interaction within demonstrations and more interaction with the last token, indicating that the model uses the accumulated task information to generate output.
@@
To observe more clear patterns, we group the layers into 3 groups: shallow (average of layers 1-14), intermediate (average of layers 15-16), and deep (layers 17 and above) in Figure˜20. We observe that $(i)$ in shallow layers, the model focuses on token-to-label mappings within demonstrations, reflecting local compression, $(ii)$ intermediate layers shift focus toward the final token, indicating integration of task-level information, and $(iii)$ deep layers emphasize interactions between the query and final token, aligning with expression and prediction. Thus, token-level interpretability also aligns with the layerwise compression-expression trajectory.
D.2 Grid-TDNV
In Section˜4.3, we find that the position of noisy demonstration affects compression. To investigate why perturbing later demonstrations results in higher TDNV at the final separation token, we conduct a more fine-grained token-level analysis by computing the layerwise TDNV for each separation token across all demonstrations. For each demonstration’s separator, we calculate the layerwise TDNV and organize them into a grid structure, referred to as the grid TDNV. As shown in Figure˜21, perturbing a demonstration at a given position most significantly increases the TDNV of the next demonstration. However, this increase gradually diminishes as more correct demonstrations are appended. This pattern suggests that the negative impact of early errors can be partially mitigated by subsequent correct examples.
Appendix E Proof of Theorem˜5.1
@@ D.2 Grid-TDNV
Appendix F Validation of the I.I.D. Assumption under Repitition
In Theorem˜5.1, we assume that each demonstration $\bm{h}_{i},i=1,\ldots,K$ is i.i.d. randomly generated from a distribution $\mathcal{H}$ on $\mathbb{R}^{d}$. To validate the importance of this assumption, we design experiments that increase the ICL length by providing additional demonstrations as input. There are two possible extension methods:
@@
As shown in Figure˜23 and Figure˜23, in distinct mode, as new demonstrations introduce novel information, the accuracy increases and TDNV decreases. However, in repeat mode, since repeated demonstrations add no new task information, the accuracy and TDNV remain largely unchanged. This proves that when the i.i.d. condition is violated, increasing the number of demonstrations does not lead to better performance and compression.
Appendix G Task vector accuracy & Early-exit accuracy
Task vector accuracy[8] refers to how accurately a LLM can perform a task using only a learned representation of the task(task vector), instead of full in-context demonstrations. As shown in Figure˜24, to evaluate the task vector accuracy, we conduct the following steps:
@@ Appendix G Task vector accuracy & Early-exit accuracy
where $_{:,K+1}$ means the $(K+1)$-th column, $\ell$ is the layer with most compression.
@@
Early-exit accuracy [9, 51] measures a model’s prediction accuracy when using intermediate-layer representations instead of the final layer for making predictions. This metric helps assess how effectively different layers encode the information needed to complete the task. Let $\bm{h}^{(\ell)}\in\mathbb{R}^{d}$ be the hidden state of the final token at layer $\ell$, and let $\mathcal{C}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{V}$ be the last-layer classifier mapping from hidden dimension $d$ to vocabulary size $V$. Then as shown in Figure˜25, the prediction is,
@@
Appendix H Task-Vector Contrastive Fine-tuning
As shown in Figure˜26, to obtain more compressed and discriminative task vectors, we add an additional contrastive loss on task vectors that explicitly encourages representation compression. Specifically, we fine-tuning a pretrained GPT-2 model on algorithmic ICL tasks domains (next letter, next two letters, previous letter, uppercase, next uppercase letter) using a combined objective as,
$\mathcal{L}=\mathcal{L}_{\mathrm{CE}}+\beta\mathcal{L}_{\mathrm{Contra}},$ (18)
where $\mathcal{L}_{\mathrm{CE}}$ is the cross-entropy loss computed only on separator tokens to predict the correct label and $\mathcal{L}_{\mathrm{Contra}}$ is a contrastive term that pulls hidden states from the same task closer while pushing those from different tasks apart:
$\mathcal{L}_{\mathrm{Contra}}=-\frac{1}{|\Omega|}\sum_{t}\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\log\frac{\exp(\mathbf{h}_{i,t}^{\top}\mathbf{h}_{j,t}/\tau)}{\sum_{\begin{subarray}{c}(a,b)\neq(i,t)\end{subarray}}\exp(\mathbf{h}_{i,t}^{\top}\mathbf{h}_{a,b}/\tau)}.$ (19)
where $\bm{h}_{i,t}$ denotes the normalized hidden state of the last token in the intermediate layer(task vector) for the $i$-th sample of task $t$, $\tau$ is the temperature, and $\Omega$ is the set of all pairs. We set $\beta=0.1$, temperature $\tau=0.07$, and batch size $100$ with equal samples from each task. Each training example provides $K=20$ context demonstrations followed by a separator token. We finetuned the model to achieve 100% ICL accuracy on all tasks.
For evaluation, we extract task vectors with a context length of $K=20$ and assess task-vector accuracy: performing zero-shot ICL with the extracted task vector injected at the same position. Higher task-vector accuracy reflects more effective task vectors. As shown in Figure˜16, task-vector contrastive fine-tuning produces better task vectors than standard fine-tuning.
To illustrate whether task-vector contrastive fine-tuning produces more compressed task vectors, we visualize hidden states from the $7$-th layer using PCA. As shown in Figure˜27, the left panel depicts task vectors extracted from model finetuned with only the cross-entropy loss, while the right panel shows vectors extracted from model finetuned with both cross-entropy and contrastive losses. The latter exhibits more distinct and tightly clustered task representations, demonstrating the effectiveness of the contrastive objective in compressing task vectors.
Figure 26: Illustration of task-vector contrastive fine-tuning.
(a) CE loss.
(b) CE loss + contrastive loss.
Figure 27: Comparison of PCA visualizations for task vectors. The task vector is extracted from models finetuned with different losses.